Computing Present Value of Terminal Period FCFF

Use the following data to compute the present value of the terminal period free cash flows to the firm for each of the four firms A, B, C, and D. The forecast horizon included four years.

A B C D
Terminal period free cash flow to the firm (FCFF) $90,450 $34,593 $119,297 $62,452
WACC 5.0% 6.2% 4.3% 11.0%
Terminal period growth rate 1.0% 1.0% 2.0% 1.5%

Round answers to the nearest whole number.

A B C D
Present value of terminal cash flows Answer Answer Answer Answer
Answer #1

Terminal cash flow for A = 90,450

WACC = 5%

Growth rate = 1%

Present value = 90,450 * (1 + 1%)/ (5% - 1%)

Present value = $2,283,862.50

Present value = $2,283,863

Terminal cash flow for B = 34,593

WACC = 6.2%

Growth rate = 1%

Present value = 34,593 * (1 + 1%)/ (6.2% - 1%)

Present value = $671,902.50

Present value = $671,903

Terminal cash flow for C = 119,297

WACC = 4.3%

Growth rate = 2%

Present value = 119,297 * (1 + 2%)/ (4.3% - 2%)

Present value = $5,290,562.61

Present value = $5,290,563

Terminal cash flow for D = 62,452

WACC = 11%

Growth rate = 1.5%

Present value = 62,452 * (1 + 1.5%)/ (11% - 1.5%)

Present value = $667,250.32

Present value = $667,250